OpenStudy (cicilybailey14):

Function rule brain teaser To rent a cabin, a resort charges $50 plus $10 per person. Part A: Write a function rule to calculate the total cost of renting the cabin. Part B: Use your rule to find the total cost for six people to stay in the cabin.

OpenStudy (anonymous):

Do you need help with the problem???

OpenStudy (cicilybailey14):

Well I think so, I did some before this but i'm not sure if there right or if I was doing it right see

OpenStudy (cicilybailey14):

and this one

OpenStudy (anonymous):

I definitely think you are on the right track

OpenStudy (anonymous):

Here is what my answer would be.....

OpenStudy (anonymous):

f(x)=50+10x f(x)=50+(10 x 6) f(x)=50+60 f(x)=??????

OpenStudy (cicilybailey14):

okay for part a? or b?

OpenStudy (cicilybailey14):

or both?

OpenStudy (anonymous):

For both, part A is asking for you to set up a function so you would be able to solve the question. Part B is asking you to solve the problem using your function from part A.....make sense???

OpenStudy (cicilybailey14):

Yes thank you so much!

OpenStudy (anonymous):

No Problem! I tried my best :)